# Can a MISL Fly? Analysis and Ingredients for Mutual Information Skill Learning

### v1

@@ 6 Experiments
The aims of our experiments(1) verifying the theoretical analysis in Sec. 4 experimentally, (2) identifying several ingredients that are key to making MISL algorithms work well more broadly, and (3) comparing our simplified algorithm CSF to prior work. Our experiments will use standard benchmarks introduced by prior work on skill learning. All experiments show means and standard deviations across ten random seeds.
@@ 6.3 Ablation Studies
We now study various design decisions of both METRA and our simplified method, aiming to identify some key factors that boost these MISL algorithms. We will conduct ablation studies on Ant again, comparing coverage of $(x,y)$ coordinates of different variants.
(1) Contrastive learning recovers METRA’s representation objective. Our analysis (Sec. 4.1) and experiments (Sec. 6.2) have shown that METRA learns contrastive representations. We now test whether we can retain the performance of METRA by simply replacing its representation objective with the contrastive lower bound (Eq. 8). Results in Fig. 3 (Left) suggest that using the contrastive loss (METRA-C) fully recovers the original performance, circumventing explanations building upon the Wasserstein dependency measure.
@@ 6.4 CSF Matches SOTA for both Exploration and Downstream performance
Figure 4: CSF performs on par with METRA. We compare CSF with baselines on state coverage (left), zero-shot goal reaching (middle), and hierarchical control (right). We generally find CSF to perform roughly on par with METRA and outperform all other baselines in most settings. Shaded areas indicate one standard deviation. Appendix Fig. 567 show the learning curves for all tasks.
@@
Experimental Setup. We evaluate on the same five tasks as those used in Park et al. (2024) plus Robobin from LEXA (Mendonca et al., 2021), though we will only focus HalfCheetah and Humanoid in the main text. For baselines, we also use a subset from Park et al. (2024) (METRA (Park et al., 2024), CIC (Laskin et al., 2022), DIAYN (Eysenbach et al., 2019), and DADS (Sharma et al., 2020)) along with VISR (Hansen et al., 2020). See Appendix B.2 for details.
Exploration performance. To measure the inherent exploration capabilities of each method without considering any particular downstream task, we compute the state coverage by counting the unique number of $(x,y)$ coordinates visited by the agent. Fig. 4 (left) shows CSF matches METRA on both HalfCheetah and Humanoid. For the full set of exploration results, please see Appendix B.3.
Zero-shot goal reaching. In this setting the agent infers the right skill given a goal without further training on the environment. We evaluate on the same set of six tasks and defer both the goal sampling and skill inference strategies to Appendix B.4. We report the staying time fraction, which is the number of time steps that the agent stays at the goal divided by the horizon length. In Fig. 4 (middle), we find all methods to perform similarly on HalfCheetah, while METRA and CSF perform best on Humanoid, with METRA performing slightly better on the latter. For the full set of zero-shot goal reaching results, please see Appendix B.4.
Hierarchical control. We train a hierarchical controller $\pi_{h}(z\mid s)$ that outputs latent skills $z$ as actions for every fixed number of time steps to maximize the discounted return in two downstream tasks from Park et al. (2024), one of which requires to reach a specified goal (HumanoidGoal) and one requires jumping over hurdles (HalfCheetahHurdle). The results in Fig. 4 (right) show CSF and METRA are the best performing methods, showing mostly similar performance. For further details as well as the full set of results on all tasks, please see Appendix B.5.
@@ 7 Conclusion
In this paper, we show how one of the current strongest unsupervised skill discovery algorithms can be understood through the lens of mutual information skill learning. Our analysis allowed the development of our new method CSF, which we showed to perform on par with METRA in most settings. More broadly, our work provides evidence that mutual information maximization can still be effective to build high performing skill discovery algorithms.
Limitations. While we find CSF to perform relatively well on the standard benchmarks (Park et al., 2024), it is unclear how far its performance can scale to increasingly complex environments such as Craftax (Matthews et al., 2024) or VIMA (Jiang et al., 2022), which present an increased number of interactive objects, partial observability, environment stochasticity, and discrete action spaces. Another open question is how to perform scalable pre-training on large datasets, e.g., BridgeData V2 (Walke et al., 2023) or YouCook2 (Zhou et al., 2018), using MISL algorithms such as CSF to get both transferable state representations and diverse skill-conditioned policies. We leave investigating these empirical scaling limits to future work.
@@ A.4 Proof of Proposition 3
$I^{\pi}(S,S^{\prime};Z)\geq h(Z)+\mathbb{E}_{p^{\pi}(s,s^{\prime},z)}[\log q^{\pi}(z\mid s,s^{\prime})],$
where $q^{\pi}(z\mid s,s^{\prime})$ is an arbitrary variational approximation of $p^{\pi}(z\mid s,s^{\prime})$. We can set $\log q^{\pi}(z\mid s,s^{\prime})$ to be
$\log q^{\pi}(z\mid s,s^{\prime})$ $=f(s,s^{\prime},z)+\log p(z)-\log\mathbb{E}_{p(z)}\left[e^{f(s,s^{\prime},z)}\right],$
@@ C.1 Quadratic approximation of $\textmd{LB}^{\beta}_{2}(\phi)$
Figure 9: Skill dimension. CSF and METRA (to a lesser extent) are sensitive to the skill dimension (indicated in parentheses).
@@ C.3 Sensitivity of CSF to the scaling coefficient $\xi$
Figure 10: CSF $\xi$ sensitivity. Increasing $\xi$ to slightly higher values (2, 5, 10) boosts performance of CSF, while a $\xi<1$ hurts performance substantially.
@@
C.4 Skill Visualizations

### v2

@@ 6 Experiments
The aims of our experiments are (1) verifying the theoretical analysis in Sec. 4 experimentally, (2) identifying several ingredients that are key to making MISL algorithms work well more broadly, and (3) comparing our simplified algorithm CSF to prior work. Our experiments will use standard benchmarks introduced by prior work on skill learning. All experiments show means and standard deviations across ten random seeds.
@@ 6.3 Ablation Studies
We now study various design decisions of both METRA and CSF, aiming to identify some key factors that boost these MISL algorithms. We will conduct ablation studies on Ant again, comparing coverage of $(x,y)$ coordinates of different variants.
(1) Contrastive learning recovers METRA’s representation objective. Our analysis (Sec. 4.1) and experiments (Sec. 6.2) have shown that METRA learns contrastive representations. We now test whether we can retain the performance of METRA by simply replacing its representation objective with the contrastive lower bound (Eq. 8). Results in Fig. 3 (Left) suggest that using the contrastive loss (METRA-C) fully recovers the original performance, circumventing the Wasserstein dependency measure.
@@ 6.4 CSF Matches SOTA for both Exploration and Downstream performance
Figure 4: CSF performs on par with METRA. We compare CSF with baselines on state coverage (left), zero-shot goal reaching (middle), and hierarchical control (right). CSF performs roughly on par with METRA and outperform all other baselines in most settings. Shaded areas indicate one standard deviation. Appendix Fig. 567 show the learning curves for all tasks.
@@
Experimental Setup. We evaluate on the same five tasks as those used in Park et al. (2024) plus Robobin from LEXA (Mendonca et al., 2021). For baselines, we also use a subset from Park et al. (2024) (METRA (Park et al., 2024), CIC (Laskin et al., 2022), DIAYN (Eysenbach et al., 2019), and DADS (Sharma et al., 2020)) along with VISR (Hansen et al., 2020). See Appendix B.2 for details.
Exploration performance. To measure the unsupervised exploration capabilities of each method, we compute the state coverage by counting the unique number of $(x,y)$ coordinates visited by the agent. Fig. 4 (left) shows CSF matches METRA on both HalfCheetah and Humanoid. See Appendix B.3 for full results on exploration.
Zero-shot goal reaching. In this setting, the agent infers the right skill given a goal without further training on the environment. We evaluate on the same set of six tasks and defer both the goal sampling and skill inference strategies to Appendix B.4. We report the staying time fraction, which is the number of time steps that the agent stays at the goal divided by the horizon length. In Fig. 4 (middle), we find all methods to perform similarly on HalfCheetah, while METRA and CSF perform best on Humanoid, with METRA performing slightly better on the latter. See Appendix B.4 for full results on zero-shot goal reaching.
Hierarchical control. We train a hierarchical controller $\pi_{h}(z\mid s)$ that outputs latent skills $z$ as actions for every fixed number of time steps to maximize the discounted return in two downstream tasks from Park et al. (2024), one of which requires to reach a specified goal (HumanoidGoal) and one requires jumping over hurdles (HalfCheetahHurdle). The results in Fig. 4 (right) show CSF and METRA are the best performing methods, showing mostly similar performance. See Appendix B.5 for full results and details.
@@ 7 Conclusion
In this paper, we show our understanding of a current SOTA unsupervised skill discovery algorithms through the lens of MISL. Our analysis allowed the development of a simpler method CSF, which performs on par with METRA in most settings. More broadly, we provide evidence that mutual information maximization can still be effective to build high performing skill discovery algorithms.
Limitations. While CSF performs relatively well on the standard benchmarks, it is unclear how to scale the performance to increasingly complex environments such as Craftax (Matthews et al., 2024) or VIMA (Jiang et al., 2022), which present an increased number of objects, partial observability, stochasticity, and discrete action spaces. Another open question is how to perform scalable pre-training on large datasets, e.g., BridgeData V2 (Walke et al., 2023) or YouCook2 (Zhou et al., 2018), using MISL algorithms such as CSF to get both transferable state representations and diverse skill-conditioned policies. We leave investigating these empirical scaling limits to future work.
Reproducibility Statement
We have made our code publicly available here: https://github.com/Princeton-RL/contrastive-successor-features. In addition, we have dedicated several sections in the appendix to further ensure reproducibility of our results. Appendix B provides a detailed account of all experimental details, including GPU types, training times, hyperparameters and architectural details (Tables 12, and 3), and detailed descriptions of the parameters of the various settings we use to compare with prior work (Appendices B.3B.4B.5). Finally, we have also included full proofs for the theoretical results stated in Proposition 12, and 3 which can be found in Appendices A.2A.3, and A.4, respectively.
@@ A.4 Proof of Proposition 3
$I^{\pi}(S,S^{\prime};Z)\geq h(Z)+\mathbb{E}_{p^{\pi}(s,s^{\prime},z)}[\log\tilde{q}(z\mid s,s^{\prime})],$
where $\tilde{q}(z\mid s,s^{\prime})$ is an arbitrary variational approximation of $p^{\pi}(z\mid s,s^{\prime})$. We can set $\log\tilde{q}(z\mid s,s^{\prime})$ to be
$\log\tilde{q}(z\mid s,s^{\prime})$ $=f(s,s^{\prime},z)+\log p(z)-\log\mathbb{E}_{p(z)}\left[e^{f(s,s^{\prime},z)}\right],$
@@ C.1 Quadratic approximation of $\textmd{LB}^{\beta}_{2}(\phi)$
Figure 9: Ablation of the skill dimension. CSF and METRA (to a lesser extent) are sensitive to the dimension of skill (indicated in parentheses).
@@ C.3 Sensitivity of CSF to the scaling coefficient $\xi$
Figure 10: The effect of $\xi$ in CSF. Increasing $\xi$ to slightly higher values (2, 5, 10) boosts performance of CSF, while a $\xi<1$ hurts performance substantially.
@@
C.4 Sensitivity of METRA to $\lambda$, $\epsilon$, and the dual learning rate
In Fig 11, we study the impact on the state coverage of METRA in Ant for various hyperparameters. Specifically, we vary the initial value of the dual variable $\lambda$, the slack variable $\epsilon$, and the learning rate of the dual variable. We find that METRA is robust to all three hyperparameters. All experiments plot the mean and standard deviation over 2 random seeds.
(a) $\lambda$ (dual variable initial value)
(b) $\epsilon$ (slack variable)
(c) Learning rate of the dual variable
Figure 11: METRA is robust to the initial value of the dual variable $\lambda$, the slack variable $\epsilon$, and the learning rate of the dual variable.
C.5 Skill Visualizations
@@
C.6 Failed Experiments
We experimented with both CSF and METRA on the MiniHack-Corridor-R5-v0 from MiniHack (Samvelyan et al., 2021). This environment has a discrete action space and requires the agent to navigate between 5 different rooms using various doors and corridors. None of these methods got consistently stable performance on this environment. We conjecture that the difficulties could come from (1) the discrete action space, (2) randomized initial states and dynamics (layout of the rooms), and (3) partial observability.